what is the quotient when you divided 2/5 by 1/10
Answer by richwmiller(17219)

multiplying by the reciprocal of the number is the same as dividing by the number
dividing by 4 is the same as multiplying by 1/4
dividing by 1/10 is the same as multiplying by 10/1 or 10
(2/5)/(1/10)=2/5*10/1=4